# VALLR: Visual ASR Language Model for Lip Reading

### v1

@@ Abstract
Lip Reading, or Visual Automatic Speech Recognition (V-ASR), is a complex task requiring the interpretation of spoken language exclusively from visual cues, primarily lip movements and facial expressions. This task is especially challenging due to the absence of auditory information and the inherent ambiguity when visually distinguishing phonemes that have overlapping visemes—where different phonemes appear identical on the lips. Current methods typically attempt to predict words or characters directly from these visual cues, but this approach frequently encounters high error rates due to coarticulation effects and viseme ambiguity. We propose a novel two-stage, phoneme-centric framework for Visual Automatic Speech Recognition (V-ASR) that addresses these longstanding challenges. First, our model predicts a compact sequence of phonemes from visual inputs using a Video Transformer with a CTC head, thereby reducing the task complexity and achieving robust speaker invariance. This phoneme output then serves as the input to a fine-tuned Large Language Model (LLM), which reconstructs coherent words and sentences by leveraging broader linguistic context. Unlike existing methods that either predict words directly—often faltering on visually similar phonemes—or rely on large-scale multimodal pre-training, our approach explicitly encodes intermediate linguistic structure while remaining highly data efficient. We demonstrate state-of-the-art performance on two challenging datasets, LRS2 and LRS3, where our method achieves significant reductions in Word Error Rate (WER) achieving a SOTA WER of 18.7 on LRS3 despite using 99.4 less labelled data than the next best approach. Code will be released following the review process.
@@ 3.3 Adapter Network with Temporal Downsampling
The sequence length of the ViT output can be prohibitively large for the subsequent CTC module. To reduce the temporal dimension, we apply a sequence of 1D convolutions and pooling layers, as shown in Table 1. Formally,
@@
Layer Name | Output Shape | Settings |
Conv1 | (B, 768, 256) | filters=256, kernel_size=5, stride=2, padding=2 |
MaxPool1 | (B, 384, 256) | pool_size=2 |
Conv2 | (B, 192, 256) | filters=256, kernel_size=3, stride=2, padding=1 |
MaxPool2 | (B, 96, 256) | pool_size=2 |
Conv3 | (B, 48, 256) | filters=256, kernel_size=3, stride=2, padding=1 |
MaxPool3 | (B, 16, 256) | pool_size=3 |
Conv4 | (B, 8, 256) | filters=256, kernel_size=3, stride=3, padding=1 |
MaxPool4 | (B, 1, 256) | pool_size=5, stride=8 |
Flatten | (B, 256) | Flatten |
@@ 7 Introduction
This supplementary material provides additional qualitative results from our proposed network architecture, evaluated on the LRS3 dataset [3]. We include video examples (referenced as exampletranslation1.mp4 and exampletranslation2.mp4) to demonstrate the effectiveness of the method in performing silent video captioning. These examples overlay the predicted captions on the original videos to offer an intuitive understanding of the predictions. Furthermore, we provide detailed analysis of common errors in phoneme-to-word reconstruction to identify limitations and strengths of the model. The additional results in this document are organized in three parts:
@@
Tab. 8: Challenging cases, including unseen names (e.g., Kofi Annan), demonstrating areas for future improvement with larger-scale pertaining.

### v2

@@ Abstract
Lip Reading, or Visual Automatic Speech Recognition (V-ASR), is a complex task requiring the interpretation of spoken language exclusively from visual cues, primarily lip movements and facial expressions. This task is especially challenging due to the absence of auditory information and the inherent ambiguity when visually distinguishing phonemes that have overlapping visemes, where different phonemes appear identical on the lips. Current methods typically attempt to predict words or characters directly from these visual cues, but this approach frequently encounters high error rates due to coarticulation effects and viseme ambiguity. We propose a novel two-stage, phoneme-centric framework for Visual Automatic Speech Recognition (V-ASR) that addresses these longstanding challenges. First, our model predicts a compact sequence of phonemes from visual inputs using a Video Transformer with a CTC head, thereby reducing the task complexity and achieving robust speaker invariance. This phoneme output then serves as the input to a fine-tuned Large Language Model (LLM), which reconstructs coherent words and sentences by leveraging broader linguistic context. Unlike existing methods that either predict words directly or rely on large-scale multimodal pre-training, our approach explicitly encodes intermediate linguistic structure while remaining highly data efficient. We demonstrate state-of-the-art performance on two challenging datasets, LRS2 and LRS3, where our method achieves significant reductions in Word Error Rate (WER) achieving a SOTA WER of 18.7 on LRS3 despite using 99.4 less labelled video data than the next best approach. Code is available here: https://github.com/MarshallT-99/VALLR
@@ 3.3 Adapter Network with Temporal Downsampling
The sequence length of the ViT output can be prohibitively large for the subsequent CTC module. To reduce the temporal dimension, we apply a sequence of 1D convolutions and pooling layers, as shown in Table LABEL:tab:adapter. Formally,
@@
| Stage | Module | In Out Channels | Kernel | Stride | Padding | Downsample Factor | Output Length |
| 1 | Conv1d | feature_size adapter_dim | 5 | 2 | 2 | ||
| BatchNorm1d | adapter_dim adapter_dim | – | – | – | |||
| ReLU | — | – | – | – | |||
| 2 | Conv1d | adapter_dim adapter_dim | 3 | 2 | 1 | ||
| BatchNorm1d | adapter_dim adapter_dim | – | – | – | |||
| ReLU | — | – | – | – | |||
| 3 | Conv1d | adapter_dim adapter_dim | 3 | 2 | 1 | ||
| BatchNorm1d | adapter_dim adapter_dim | – | – | – | |||
| ReLU | — | – | – | – | |||
| 4 | Conv1d | adapter_dim adapter_dim | 3 | 3 | 1 | ||
| BatchNorm1d | adapter_dim adapter_dim | – | – | – | |||
| ReLU | — | – | – | – | |||
| 5 | AvgPool1d | adapter_dim adapter_dim | 5 | 8 | 0 | ||
| 6 | Linear | adapter_dim CTC_dim | - | - | - | ||
| ReLU | — | – | – | – |
@@ 6 Conclusion and Limitations
Acknowledgements
This work was supported by the SNSF project ‘SMILE II’ (CRSII5 193686), the Innosuisse IICT Flagship (PFFS-21-47), EPSRC grant APP24554 (SignGPT-EP/Z535370/1) and through funding from Google.org via the AI for Global Goals scheme. This work reflects only the author’s views and the funders are not responsible for any use that may be made of the information it contains.
@@ 7 Introduction
This supplementary material provides additional qualitative results from our proposed network architecture, evaluated on the LRS3 dataset [3]. We include video examples referenced as exampletranslation1.mp4 and exampletranslation2.mp4 to demonstrate the effectiveness of the method in performing silent video captioning. These examples overlay the predicted captions on the original videos to offer an intuitive understanding of the predictions. Furthermore, we provide detailed analysis of common errors in phoneme-to-word reconstruction to identify limitations and strengths of the model. The additional results in this document are organized in three parts:
@@
Tab. 8: Challenging cases, including unseen names (e.g., Kofi Annan), demonstrating areas for future improvement with larger-scale pretraining.